# Can GenAI Move from Individual Use to Collaborative Work? Experiences, Challenges, and Opportunities of Integrating GenAI into Collaborative Newsroom Routines

### v2

@@ 7.4. Creating New Work Routines around GenAI
Several participants proposed creating entirely new work routines that are uniquely enabled by GenAI, workflows that would be difficult, if not impossible, to realize under traditional human-only newswork. These routines do not simply replicate existing editorial practices with AI assistance; rather, they reflect a radical transformation in how content is produced, how collaboration unfolds, and how audiences engage with journalistic output in the age of GenAI.
@@
Firstly, a common suggestion was the creation of dedicated AI news zones within newsrooms: editorial spaces deliberately separated from core news operations, designed to foster experimentation with GenAI tools. As P2 put it, “In these AI news zones, people know it’s okay to try something new. We’re not pretending it’s traditional reporting. It’s a different GenAI product.”
@@
These AI zones were described as low-risk, high-flexibility environments where journalists teams could develop novel formats and experiment with creative applications of GenAI. Participants envisioned a wide range of possibilities that would be difficult to achieve through conventional human-only news production. For instance, some proposed GenAI-powered chatbots that “allow audiences to converse with digital journalist avatars to better understand complex news stories” (P22). Short-form videos and animations, scripted with GenAI and targeted at younger, mobile-first audiences, were also frequently mentioned, with P7 noting that “there’s already a lot on Douyin, and they’re popular”. Still others proposed AIGC documentaries composed from archival footage and GenAI-generated narration. P11 explained: “It’s technically feasible, and we have extensive archives, especially rich local footage that can be used to create documentaries about local cultures.”
@@
Because these GenAI products are clearly distinct in form and tone from conventional journalism, our participants believe audiences tend to be more tolerant of stylistic experimentation and minor inaccuracies. Editors, in turn, feel freer to iterate on editorial norms. As P11 reflected, “When people see AI-generated content in these special AI news zones, they don’t judge it heavily by the standards of our front page. They expect it to be more innovative.”
@@
Opening new routines for GenAI does not eliminate editorial control in our participants’ expectations. As P8 emphasized, even content produced in AI news zones must undergo human review, be clearly labeled as AI-generated, and be documented internally. However, the emergence of these formats offers a structured space to explore the creative frontier of GenAI without forcing immediate reform of core editorial systems. “In these zones,” P8 added, “journalists and editors can be more creative in how they collaborate with GenAI, without fear of professional judgment or reputational risk.”
@@
Secondly, several participants also envisioned a new collaborative newswork form augmented by GenAI. P2 highlighted an experimental system called Fabarta Agents (E-commerce Research Center of China (2025), NetEase E-Commerce), currently being trialed at CCTV and several provincial news televisions as an internal newsroom application. A key feature of the GenAI system is its deliberate emphasis on peer learning and stylistic imitation. Journalists can recall the full corpus of a preferred journalist colleague’s past work in the system, using it not only as training material but also as real-time reference during writing. The agent can, for example, suggest opening paragraphs, headlines, or structural outlines based on how a trusted peer might have approached similar stories. As P2 explained, “You can ask the agent to write in the style of someone you respect in the newsroom. It’s like getting mentorship through the GenAI system.”
@@
According to P2, Fabarta Agents opens up a new pathway for collegial influence and inspiration, where AI becomes a medium for professional emulation and reflection, not merely automation. As P2 summarized, “To me, this GenAI tool essentially creates a safe space for experimentation while also improving our collaborative culture. It shows how technology itself can reshape the very nature of collaboration.”
@@ 8. Discussion
Our findings suggest that while journalists widely adopt GenAI tools at the individual level, structural and cultural barriers prevent these practices from scaling into sustained team-based routines. At the same time, journalists envision productive futures for GenAI-augmented collaboration. These findings offer important design implications for HCI and CSCW researchers seeking to embed GenAI into collaborative work, particularly in high-stakes, institutionalized domains such as journalism.
This discussion section unpacks three interrelated dimensions. First, it calls for HCI and CSCW research to move beyond a predominant focus on individual interactions with GenAI and instead attend to its role in collaborative newsroom work (subsection 8.1). Second, it rethinks GenAI’s distinctive technical essence in contemporary collaboration, not as a peripheral support tool but as a system that directly intervenes in the epistemic and creative core of journalism, thereby reshaping how collaboration itself must be organized and legitimated (subsection 8.2). Third, it outlines concrete design and policy implications to support more transparent, accountable, and responsible forms of GenAI-augmented collaboration (subsection 8.3).
@@ 8.1. From Individual GenAI Use to Collaborative Work
While recent HCI and CSCW literature has explored the growing role of GenAI in creative and knowledge work, most of this research has focused on individual interactions with AI systems: how users prompt, evaluate, and iteratively shape GenAI outputs to support tasks such as writing, ideation, or coding (Palani and Ramos, 2024; Sun et al., 2024; Lee et al., 2025; Li et al., 2024b). In contrast, our study contributes to an emerging but underdeveloped area: how GenAI is (or is not) integrated into collaborative work practices, particularly in high-stakes, institutionally-bound domains such as journalism. It also raises forward-looking questions about how organizations can embed GenAI into future collaborative routines.
Our findings reveal that GenAI-augmented collaboration is not a natural extension of individual use, but a site of organizational tension, negotiation, and technical infrastructural fragility. Despite widespread individual experimentation with GenAI usage, newsrooms struggle to transform these fragmented practices into shared routines or team-level workflows. This aligns with long-standing HCI and CSCW insights that new technologies rarely slot in to existing work practices without reconfiguration of roles, norms, and communication structures (Møller, 2024; Holtgrewe, 2014; Blomberg and Karasti, 2013; Geiger et al., 2021; Hu et al., 2022). To some extent, adopting new technology also means adopting potential cultural or organizational transformations along with technology adoption. GenAI collaboration in our cases is a socio-technical challenge, encompassing both the design of collaborative interfaces and the need to ensure accountability, visibility, interpretability, and trust among newsroom actors.
In the technical aspect, our study suggests that the opacity of GenAI interactions, both in how prompts are crafted by journalists and in how outputs are generated by GenAI, renders shared understanding and joint decision-making difficult. This dual black-boxing of human prompting and machine generation creates asymmetries in editor–journalist collaboration, echoing prior HCI and CSCW concerns about collaboration breakdowns in common ground when collaborators cannot access or reconstruct one another’s interactions with complex systems (Carroll et al., 2006; Fischer, 2004; Ackerman et al., 2013). Editors, in our study, expressed a reluctance to incorporate GenAI into collaborative routines precisely because they lack technical mechanisms for auditability, revision tracking, or interpretive control.
@@
While prior HCI studies have explored the technical design of AI tools to support individual workflows (Lu et al., 2025; Zhu et al., 2024), and CSCW scholars have examined the breakdown of collaboration in opaque systems (Easterbrook, 2012; Gui et al., 2018), our findings foreground a distinct dynamic: GenAI use becomes routinized individually without becoming institutionalized in collaborative work. This decoupling of individual technical use and limited social legitimacy, of what people do versus what they are willing to show, poses a unique barrier to scaling GenAI from individual to collaborative work.
While GenAI is increasingly adopted across real-world organizations, it has not yet emerged as an immediate or evident enabler of collaborative intelligence (Rajaram and Tinguely, 2024; Perifanou and Economides, 2025). This calls for further exploration into how GenAI can be more effectively integrated into organizational practices. In practice, moving from individual to collective GenAI use requires deep institutional scaffolding: shared norms about appropriate use, technical infrastructures that enable co-editing, and cultural shifts around disclosure and responsibility. We argue that GenAI does not “scale” to collaboration by default: it must be domesticated into organizational life through deliberate design and negotiation.
8.2. Rethinking GenAI as a Collaborative Technology in Future Newsroom
@@
Unlike earlier collaborative technologies that operated behind the scenes, GenAI directly intervenes in the epistemic and creative core of journalism. This shift forces a renegotiation of routines, roles, and professional legitimacy. Prior HCI, CSCW and organizational research has shown how emerging technologies, such as groupware, workflow systems, and enterprise platforms, reshape collaboration by redistributing tasks, restructuring coordination, and altering authority (Rama and Bishop, 2006; Poltrock and Grudin, 1998; Grudin, 1999). These previous generations of tools, however, typically operated in the background: they supported communication, scheduling, or documentation, but rarely intervened in the substantive outputs of professional work (Rama and Bishop, 2006; Knearem et al., 2023; Xiao et al., 2025a). Consequently, their adoption was often gradual and partial, limited to situations where technology could be integrated without disrupting established routines of teamwork and accountability.
GenAI differs in a fundamental way. Rather than remaining in the background, it intervenes directly and radically in the epistemic and creative core of journalism by drafting headlines, summarizing interviews, or generating full article drafts, as we mentioned in subsection 6.2 and subsection 7.2. This direct involvement not only reshapes how work is done but also blurs professional boundaries of authorship, editorial authority, and responsibility. In contrast to earlier technologies that modified supporting processes, GenAI compels today’s journalists and editors to renegotiate what counts as legitimate professional newswork and who is accountable for it.
We argue that GenAI’s intervention in core editorial tasks fundamentally alters how collaboration unfolds in the newsroom. We found that journalists rarely disclosed their practices to colleagues or editors. This secrecy was driven not only by coordination difficulties, such as tracing AI contributions, but also by deeper anxieties. Many feared that reliance on GenAI would erode their professional skills, increase the risk of replacement, or compromise copyright and recognition of their work, echoing earlier HCI research on copyright concerns (Suresh et al., 2024) and secret GenAI use (Zhang et al., 2025a).
These findings highlight that, for GenAI to support rather than disrupt newsroom collaboration, it must move beyond hidden, individualized use, and become embedded in shared routines. This direct intervention from GenAI in the substance of journalistic work transforms collaboration at a fundamental level: questions of authorship, authority, and accountability are no longer peripheral adoption issues but central to everyday teamwork in the age of GenAI. Secretive reliance on GenAI, as discussed in subsection 5.2 and subsection 6.2, can by contrast actively undermine collaboration, eroding trust among colleagues and obscuring accountability for editorial decisions. These practices often stem from journalists’ desire to preserve their professional identity and safeguard the distinct value of human judgment, even as GenAI encroaches on the very tasks that define their role. Yet such consequences underscore the need to fundamentally reconsider GenAI as a technology that disrupts the epistemic and creative core of journalism, and to reimagine how collaborative practices can adapt in response.
We thus urge journalism scholars and practitioners to move beyond treating GenAI as a matter of private experimentation. Instead, attention should turn to how newsroom cultures can evolve to support more transparent, coordinated, and principled uses of GenAI that reinforce journalism’s collective standards of fairness, independence, and accuracy. Future HCI and CSCW research should investigate how GenAI’s integration into collaborative routines reshapes professional identity, and how, in this context, teams and organizations can renegotiate the boundaries of human judgment and GenAI contribution in sustaining the professional identity. More broadly, scholars must also consider how a technology as disruptive as GenAI can be embedded into future collaboration in ways that channel its interventions into collective, accountable, and sustainable practices rather than isolated or hidden uses.

### v3

@@ 7.4. Creating New Workflow Around GenAI
Several participants proposed creating entirely new workflow that are uniquely enabled by GenAI, workflows that would be difficult, if not impossible, to realize under traditional human-only newswork. These workflow shall reflect a transformation in how content is produced, how collaboration unfolds, and how audiences engage with journalistic output in the age of GenAI.
7.4.1. Firstly, a common suggestion was the creation of dedicated AI news zones within newsrooms.
@@
These editorial spaces are deliberately separated from core news operations and designed to foster experimentation with GenAI tools. As $P2_{\text{female, editor/manager}}$ put it, “In these AI news zones, audiences know it’s okay to try something new. We’re not pretending it’s traditional reporting.”
@@
These AI zones were described as low-risk, high-flexibility environments where journalists teams could develop novel formats and experiment with creative applications of GenAI. Participants envisioned a wide range of possibilities that would be difficult to achieve through conventional human-only news production. For instance, some proposed GenAI-powered chatbots that “allow audiences to converse with digital journalist avatars to better understand news stories” ($P22_{\text{male, journalist}}$). Short-form videos and animations, scripted with GenAI and targeted at younger, mobile-first audiences, were also frequently mentioned, with $P7_{\text{male, editor}}$ noting that “there’s already a lot on Douyin, and they’re popular”. Still others proposed AIGC documentaries composed from archival footage and GenAI-generated narration. $P11_{\text{female, journalist}}$ explained: “It’s technically feasible, and we have extensive archives, especially rich local footage that can be used to create documentaries about local cultures.”
@@
Because these GenAI products are clearly distinct in form and tone from conventional journalism, our participants believe audiences tend to be more tolerant of stylistic experimentation and minor inaccuracies. Editors, in turn, feel freer to iterate on editorial norms. As $P11_{\text{female, journalist}}$ reflected, “When people see AI-generated content in these special AI news zones, they don’t judge it heavily by the standards of our front page. They expect it to be more innovative.”
@@
Opening new workflow for GenAI does not eliminate editorial control in our participants’ expectations. As $P8_{\text{male, editor}}$ emphasized, even content produced in AI news zones could undergo human review, be clearly labeled as AI-generated, and be documented internally. However, the emergence of these formats offers a structured space to explore the creative frontier of GenAI without forcing immediate reform of core editorial systems. “In these zones,” $P8_{\text{male, editor}}$ added, “journalists and editors can be more creative in how they collaborate with GenAI, without fear of professional judgment or reputational risk.”
Despite this enthusiasm for clearly demarcated AI news zones, many participants cautioned that even explicit GenAI labeling could become problematic when AI-generated content appears outside these designated spaces. Several participants (such as $P11_{\text{female, journalist}}$ and $P24_{\text{female, journalist}}$) worried that if a GenAI-labeled product is placed alongside traditional news stories, the label itself may trigger suspicion among readers and colleagues. Rather than reassuring audiences, an AI tag in a conventional news section could lead them to question the authenticity of surrounding content, prompting concerns such as “If this one is AI-generated, how many others might be as well?” ($P24_{\text{female, journalist}}$) Participants emphasized that, without the protective boundary of AI zones, GenAI labels risk contaminating the credibility of adjacent human-written stories and creating unnecessary verification work. As $P14_{\text{female, editor/journalist}}$ said, “Even if you label it, once AI shows up in the regular news pages, people start doubting everything around it.” For this reason, journalists viewed spatial and categorical separation, not labeling alone, as essential to preserving trust, clarity, and workflow stability in the broader newsroom environment.
7.4.2. Secondly, several participants also envisioned a new collaborative newswork form augmented by GenAI.
@@
$P2_{\text{female, editor/manager}}$ highlighted an experimental system called Fabarta Agents (E-commerce Research Center of China (NetEase E-Commerce), 2025), currently being trialed at CCTV and several provincial news televisions as an internal newsroom application. A key feature of the GenAI system is its deliberate emphasis on peer learning and stylistic imitation. Journalists can recall the full corpus of a preferred journalist colleague’s past work in the system, using it not only as training material but also as real-time reference during writing. The agent can, for example, suggest opening paragraphs, headlines, or structural outlines based on how a trusted peer might have approached similar stories. As $P2_{\text{female, editor/manager}}$ explained, “You can ask the agent to write in the style of someone you respect in the newsroom. It’s like getting mentorship through the GenAI system.”
@@
According to $P2_{\text{female, editor/manager}}$, Fabarta Agents opens up a new pathway for collegial influence and inspiration, where AI becomes a medium for professional emulation and reflection, not merely automation. As $P2_{\text{female, editor/manager}}$ summarized, “To me, this GenAI tool essentially creates a safe space for experimentation while also improving our collaborative culture. It shows how technology itself can reshape the very nature of collaboration.” For $P8_{\text{male, editor}}$, one major challenge in developing internal AI tools is “copyright”. He noted that “journalists are often worried about how their data might be used or repurposed by the system”. In his view, a well-designed workflow could help alleviate these concerns by clearly establishing boundaries on how content is stored and reused, and ensuring that individual contributions are protected.
@@ 8. Discussion
Our findings suggest that while journalists widely adopt GenAI tools at the individual level, structural and cultural barriers prevent these practices from scaling into sustained teamwork. At the same time, journalists envision productive futures for GenAI in collaboration. These findings offer important design implications for HCI and CSCW researchers seeking to embed GenAI into collaborative work, particularly in high-stakes, institutionalized domains such as journalism.
This discussion section unpacks three interrelated dimensions. First, it calls for HCI and CSCW research to move beyond a predominant focus on individual interactions with GenAI and instead attend to its role in collaborative newsroom work (subsection 8.1). Second, it rethinks GenAI’s technical essence in contemporary collaboration, as GenAI could help with writing and editing or other creative tasks, thereby affecting how collaboration itself could be organized and legitimated (subsection 8.2). Thirdly, drawing from the Chinese newsroom context, we extend these insights to broader cultural and institutional settings (subsection 8.3). Finally, it outlines concrete design and policy implications to support more transparent, accountable, and responsible forms of future coordination of GenAI into collaborative newswork (subsection 8.4).
@@ 8.1. From Individual GenAI Use to Collaborative Work
While recent HCI and CSCW literature has explored the growing role of GenAI in creative and knowledge work, most of this research has focused on individual interactions with AI systems: how users prompt, evaluate, and iteratively shape GenAI outputs to support tasks such as writing, ideation, or coding (Palani and Ramos, 2024; Sun et al., 2024; Lee et al., 2025; Li et al., 2024; Ma et al., 2025). In contrast, our study contributes to an emerging but underdeveloped area: how individualized GenAI use intersects with collaborative work, particularly in high-stakes, institutionally-bound domains such as journalism. It also raises forward-looking questions about how organizations can embed GenAI into future collaboration.
Our findings reveal that despite widespread individual experimentation with GenAI usage, newsrooms struggle to transform these fragmented practices into shared norms or team-level workflows. This aligns with long-standing HCI and CSCW insights that new technologies rarely slot in to existing work practices without reconfiguration of roles, norms, and communication structures (Møller, 2024; Holtgrewe, 2014; Blomberg and Karasti, 2013; Geiger et al., 2021; Hu et al., 2022; Ma et al., 2024). Coordinating GenAI into collaboration in our cases is a socio-technical challenge, encompassing both the design of collaborative interfaces and the need to ensure accountability, visibility, interpretability, and trust among newsroom actors.
In the technical aspect, our study suggests that the opacity of GenAI interactions—both in how prompts are crafted by journalists and in how outputs are generated by the model—renders shared understanding and joint decision-making difficult, a dynamic that participants referred to as “dual black-boxing.” This dual black-boxing of human prompting and machine generation creates asymmetries in editor–journalist collaboration, echoing prior HCI and CSCW concerns about collaboration breakdowns in common ground when collaborators cannot access or reconstruct one another’s interactions with complex systems (Carroll et al., 2006; Ackerman et al., 2013).
@@
The challenges and opportunities of moving from individualized GenAI use to collaborative newsroom practices also hinge on how different roles perceive the risks and benefits of coordination. For front-line journalists, GenAI provides pragmatic relief under heavy workloads, yet coordination also risks exposing invisible writing labor to new forms of evaluation. Their ambivalence reflects a tension between the practical benefits of GenAI and the reputational concerns associated with making its use visible. Editors encounter the dual black box from the opposite direction: without access to either prompting processes or model reasoning, they face heightened uncertainty in verification and accountability. Although shared standards could reduce ambiguity and improve review processes, editors also worry that formal coordination might institutionalize practices that could undermine editorial authority or public credibility. Managers approach GenAI from an institutional perspective shaped by efficiency, compliance, and organizational reputation. While coordinated GenAI systems are often envisioned as potential infrastructures for improving workflow consistency, managers remain wary of the political and ethical risks that organizational endorsement could entail. Informal use localizes risk to individual journalists, whereas coordinated use redistributes liability upward, increasing exposure at the organizational level. Together, these differentiated stances explain why individualized GenAI use persists even amid calls for coordination: the move from private experimentation to collective practice entails uneven risk redistribution across roles, making coordination a site of negotiation rather than a straightforward technical transition.
8.2. GenAI as a Collaborative Technology in Future Newsroom
@@
Unlike earlier collaborative technologies that supported newsroom work primarily in peripheral ways, GenAI is entering news production at points much closer to the substantive core of journalistic tasks. Rather than replacing existing collaborative norms outright, this shift introduces new pressures for renegotiating roles, workflows, and professional legitimacy. Prior HCI, CSCW, and organizational research has long demonstrated that technologies such as groupware, workflow systems, and enterprise platforms reshape collaboration by redistributing tasks, restructuring coordination, and altering authority (Rama and Bishop, 2006; Poltrock and Grudin, 1998; Grudin, 1999). These earlier systems, however, generally operated around the margins of knowledge work: they facilitated communication, scheduling, or documentation but rarely engaged directly with the content that defined professional outputs (Rama and Bishop, 2006; Knearem et al., 2023; Xiao et al., 2025b).
GenAI introduces similar dynamics of negotiation, yet with a different locus of impact. Instead of remaining in the background, GenAI more directly interacts with the content of journalistic work by drafting headlines, summarizing interviews, or generating preliminary article versions, as discussed in subsection 6.2 and subsection 7.2. This integration into content creation surface tensions more immediately tied to authorship, editorial authority, and responsibility. In this sense, GenAI prompts journalists and editors to articulate what counts as legitimate professional practice and how accountability should be shared.
Our findings find that journalists rarely disclosed their use of GenAI to colleagues or editors, not only due to coordination challenges such as tracing AI contributions, but also because of concerns about skill erosion, replacement, and copyright—all themes that resonate with prior HCI work on participation and secret AI use (Suresh et al., 2024; Zhang et al., 2025; Waters-Lynch et al., 2025). These anxieties illustrate that GenAI’s integration into newsroom collaboration intersects with existing professional norms and identities, rather than acting as a singular, disruptive force.
For GenAI to contribute constructively to collaborative newswork, practices may need to evolve beyond hidden, individualized use toward more visible and shared organizational norms. As shown in subsection 5.2 and subsection 6.2, secrecy can undermine trust, obscure accountability, and introduce uncertainty into team-based decision-making. These dynamics arise because its current use intersects with long-standing professional expectations about originality, expertise, and editorial judgment. Thus, GenAI draws attention to fundamental questions about how journalistic work is defined and evaluated, and how collaborative practices might adapt when automated tools participate more directly in core writing tasks.
Journalism scholars and practitioners may therefore benefit from shifting attention from private experimentation toward the cultural and organizational conditions that enable more transparent and principled uses of GenAI. Future HCI and CSCW research could examine how the introduction of GenAI into collaborative newswork reframes professional identity, and how teams and organizations negotiate boundaries between human judgment and computational assistance. More broadly, embedding GenAI into collaborative practice involves the familiar challenge of aligning new technological capabilities with collective standards of fairness, independence, and accuracy, rather than relying on isolated or concealed forms of use.
In many ways, GenAI extends long-standing themes in CSCW and organizational studies: new technologies enter existing ecologies of work, redistribute tasks, and prompt negotiations over authority, coordination, and accountability (Valentine et al., 2017; Xiao et al., 2025c; Herrmann and Pfeiffer, 2023). As with earlier collaborative systems, GenAI’s adoption unfolds through situated practice, shaped by organizational history, local norms, and the social meanings attached to technological use. At the same time, GenAI differs in where these negotiations occur. Whereas prior systems primarily reshaped the supporting infrastructures of collaboration, GenAI operates at the level of content creation itself, making questions of authorship, voice, and professional judgment more salient (Lewis et al., 2025; Munoriyarwa and de-Lima-Santos, 2025). This shift reorientate familiar CSCW concerns toward domains of work that were previously less technologically mediated. In other words, GenAI follows established patterns of how collaborative technologies become domesticated in organizations, yet it situates these dynamics closer to journalistic labor, where professional identity and accountability are most tightly held in writing, editing and creativity.
8.3. Extending the Chinese Insights to Other Cultural and Institutional Contexts
While our study is situated in the Chinese newsroom context, the dynamics we identify extend beyond China and offer broader implications for theorizing GenAI-supported collaboration across diverse media systems. Our findings do not merely describe a local phenomenon; instead, they surface core socio-technical patterns that can inform how media organizations in other countries understand and govern GenAI in collaborative newswork, echoing prior research on China-based cases that has generated insights applicable to global contexts (Xiao et al., 2025a; Liu, 2023; Kuai et al., 2022).
First, China provides a uniquely revealing site for studying GenAI integration because newsrooms operate under strong institutional directives to demonstrate technological modernization and political responsiveness (Xiao et al., 2025c; Kuai, 2024; Kuai et al., 2025). Participants described how AI system-building was often oriented toward external legitimacy, for instance, satisfying government expectations for digital transformation, rather than addressing the everyday collaborative practices of journalism. This top-down pressure creates symbolic, rather than operational, GenAI integration, generating a misalignment between organizational incentives and newsroom-level collaborative needs. Our findings therefore highlight a mechanism: when GenAI adoption is driven by external legitimacy pressures rather than newsroom collaborative needs, collaboration may suffers. Although political oversight varies across countries, similar tensions may emerge in media systems influenced by market competition, corporate mandates, or public scrutiny. This suggests that our Chinese case offers a conceptual lens to diagnose how external institutional pressures shape the collaborative form of GenAI coordination in other contexts.
Second, China’s hierarchical newsroom structures reveal how GenAI becomes entangled with power asymmetries in collaborative work. Editors and managers bear political and reputational risk, making them more cautious about GenAI-assisted workflows, while junior journalists privately use GenAI to cope with workload demands (Kuai, 2025). This pattern reflects a broader insight: GenAI adoption is fundamentally shaped by local distributions of authority, responsibility, and risk. Even in less centralized media systems, power asymmetries, between senior and junior journalists, editors and freelancers, or center and periphery newsrooms, can strongly influence who adopts GenAI, who discloses its use, and who bears responsibility for errors. Our findings therefore offer a transferable framework for examining how GenAI adoption interacts with organizational hierarchies across diverse professional cultures.
Third, China’s cultural norms around experimentation, public accountability, and media professionalism shape how journalists disclose or conceal GenAI use. In our study, the backstage nature of GenAI adoption reflects not only coordination challenges but also deeper cultural anxieties about maintaining credibility in a tightly regulated informational environment. Journalists feared that admitting GenAI use might be read as violating professional identity, weakening editorial rigor, or risking political misinterpretation. These findings extend Goffman’s frontstage/backstage framework (Goffman, 1949) by illustrating how backstage practices are shaped not only by personal identity work but also by broader sociopolitical conditions. This insight is generative for cross-cultural studies: transparency and disclosure norms around GenAI are likely to vary globally and will co-evolve with local journalistic values, professional identities, and cultural expectations about automation.
These insights show that the Chinese case is a theoretically productive site that foregrounds a triad of forces, i.e., political oversight, organizational hierarchy, and cultural norms, that influence GenAI’s role in collaboration. These forces are present to varying degrees across other media ecosystems. The Chinese case thereby provides both analytical tools and empirical reference points that other researchers can adapt when investigating GenAI adoption in their own institutional and cultural contexts.